OpenStudy (anonymous):

can you tell me 5x+89=79+7x plz tell the value of x

OpenStudy (anonymous):

2x=10 x=5

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

giv me medal yar

OpenStudy (anonymous):

give me medal plz

OpenStudy (anonymous):

de diys maedal sai

OpenStudy (anonymous):

tu bhi marko de de plz

OpenStudy (anonymous):

5x+89=79+7x -7x+5x= -89+79 -2x= -10 x=10/2 x=5